Circular No. 6335

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) M. Senay, University of Hawaii (UH); H. Matthews, Joint Astronomy

Centre, Hilo, and Herzberg Institute of Astrophysics, Ottawa; and D. Jewitt, UH, write: “We report the detection of CO (3-2) emission from C/1996 B2 at the James Clerk Maxwell Telescope on Mauna Kea on Mar. 1.54 and 2.52 UT. Observations on Mar. 1.54 showed the line at −57.96 ± 0.03 km/s, with peak antenna temperature 0.32 K, area 0.37 ± 0.02 K km/s, width 1.08 ± 0.07 km/s, and an ephemeris geocentric radial velocity −57.94 km/s. All quantities were confirmed on Mar. 2.52. Assuming rotational temperatures of 30 and 50 K, we estimate log Q(CO) = 28.00 and 28.04, respectively. This compares with log Q(OH) = 28.73 on Mar. 1.5 (IAUC 6333). The implied Q(CO)/Q(OH) = 0.2 is similar to CO/H_{2}O = 0.15 derived for 1P/Halley (Woods et al. 1986, Nature 324, 436). Available evidence thus points to a comet that is like 1P/Halley both in its absolute gas production and in its CO/OH and HCN/OH (IAUC 6318) ratios. Sample spectra may be exam-

ined at http://www.ifa.hawaii.edu/faculty/jewitt/submm.html.”

GRO J1744–28 B. A. Harmon, Marshall Space Flight Center, NASA; J. van Paradijs

and W. S. Paciesas, University of Alabama, Huntsville; and S. N. Zhang and C. Kouveliotou, Universities Space Research Association, report for the BATSE Team: “The 20- to 100-keV flux of GRO J1744–28, as measured with BATSE using earth occultations, has steadily decreased from 0.9 photon cm^{−2} s^{−1} on Feb. 19 to 0.5 photon cm^{−2} s^{−1} on Mar. 2. This may signal the upcoming end of the outburst of this unique source. No convincing optical/infrared counterpart of the source has so far been found, likely due to its high interstellar extinction (Augusteijn et al., IAUC 6326). The non-detection of a V -band counterpart implies that the interstellar column density, N_{H}, is at least 2 × 10^{22} cm^{−2}. At a source distance of 7 kpc, the unabsorbed apparent K magnitude is expected to be ∼ 13. Since the K-band extinction, A(K), increases with column density by A(K)/N_{H} ∼ 7.5 × 10^{−23} mag cm^{2}, the only way to identify a counterpart would appear to be deep K-band imaging centered on the variable radio source (IAUC 6323) down to at least mag 19. In view of the possible upcoming termination of the outburst, we urge observers that such images be made as soon as possible.”

1996 March 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green